# Deformation of Fuchsian representations and proper affine actions

### v1

@@ 2.1. Representation of Lie groups
We choose a connected component of $\mathfrak{a}\setminus\cup_{\alpha\in\Sigma}\ker(\alpha)$ and denote it by $\mathfrak{a}^{++}$. The closure of $\mathfrak{a}^{++}$ is denoted by $\mathfrak{a}^{+}$. Let $\mathsf{K}\subset\mathsf{G}$ (respectively $\mathsf{A}\subset\mathsf{G}$) be the connected subgroup whose Lie algebra is $\mathfrak{k}$ (respectively $\mathfrak{a}$) and let $\mathsf{A}^{+}:=\exp{(\mathfrak{a}^{+})}$. The Lie group $\mathsf{K}$ is a maximal compact subgroup of $\mathsf{G}$. The subset of restricted roots which take positive values on $\mathfrak{a}^{+}$ is denoted by $\Sigma^{+}$. We note that $\Sigma=\Sigma^{+}\sqcup-\Sigma^{+}$ and define
@@
In order to guarantee existence of proper affine actions, henceforth, we only consider representations, $\mathtt{R}$, for which $\mathsf{V}^{0}$ is nontrivial. Moreover, we assume that $\mathtt{R}$ admits an extreme, symmetric and generic vector and we fix one such vector $X_{0}$ in $\mathfrak{a}^{+}$. Such a representation $\mathtt{R}$ is called a non-swinging representation.
@@
Let $\Sigma^{+}(X_{0}):=\{\alpha\mid\alpha(X_{0})\geqslant 0\}$. We denote the vector subspace generated by $\mathfrak{g}^{0}$ and all $\mathfrak{g}^{\alpha}$ such that $\alpha\in\pm\Sigma^{+}(X_{0})$ by $\mathfrak{p}^{\pm,0}$. Let $g\in\mathsf{G}$. We say $g$ is elliptic if $g\in\mathsf{K}$. We say $g$ is hyperbolic if $g\in\mathsf{A}$. We say $g$ is unipotent if $g\in\mathsf{N}$.
@@ 2.2. Anosov representations
The pair $(\mathsf{V}^{+,0},\mathsf{V}^{-,0})$ is a transverse pair of parabolic subspaces and for any pair of transverse parabolic spaces $(\mathsf{V}_{i},\mathsf{V}_{j})$, there exists $h_{i,j}\in\mathsf{G}$ such that $(\mathsf{V}_{i},\mathsf{V}_{j})=h_{i,j}(\mathsf{V}^{+,0},\mathsf{V}^{-,0})$. If $\mathsf{V}_{i}$, $\mathsf{V}_{j}$ and $\mathsf{V}_{k}$ are three parabolic subspaces which are mutually transverse to each other and $h_{i,j},h_{i,k},h_{k,j}\in\mathsf{G}$ be such that
@@ 3.1. Neutral and neutralized sections
where the fiber over $p\in\widetilde{\mathsf{U}_{0}\Gamma}$ is denoted by $(p,\mathsf{V}^{+}_{p_{+}},\mathsf{V}^{0}_{p},\mathsf{V}^{-}_{p_{-}})$. We observe that $(p,\mathsf{V}^{+}_{p_{+}},\mathsf{V}^{0}_{p},\mathsf{V}^{-}_{p_{-}})=(p,g_{p}\mathsf{V}^{+},g_{p}\mathsf{V}^{0},g_{p}\mathsf{V}^{-})$ and $\mathsf{V}^{\pm}_{p_{\pm}}\oplus\mathsf{V}^{0}_{p}=\mathsf{V}^{\pm,0}_{p_{\pm}}$. We observe that for any $Y\in\mathsf{V}^{0}$ we have $\nu_{Y}(hg)=h\nu_{Y}(g)$ for all $g,h\in\mathsf{G}$. Hence, it follows that $\rho(\gamma)\nu_{Y}\circ\xi(\gamma_{+},\gamma_{-})=\nu_{Y}\circ\xi(\gamma_{+},\gamma_{-})$. We abuse notation and denote $\nu_{Y}\circ\xi$ by $\nu_{Y}$. These maps induce the following sections which we again denote using the same notation:
@@ 3.2. Partial affine Anosov representations
We observe that $(g,Y)\mathsf{A}_{g,Y}=\mathsf{A}_{g,Y}$. Suppose $\mathsf{A}_{1}$ and $\mathsf{A}_{2}$ are two such subspaces which are parallel to $\mathsf{V}^{0}_{g}$. Then $\mathsf{A}_{1}=\mathsf{A}_{2}+Z$ for some $Z\in\mathsf{V}$ and $gZ=Z$. Therefore, $Z\in\mathsf{V}^{0}_{g}$ and it implies that $\mathsf{A}_{1}=\mathsf{A}_{2}+Z=\mathsf{A}_{2}$, proving uniqueness. ∎
@@ 4.2. Margulis invariants
Suppose $X_{\pm},X,X_{g,Y}$ are any four points respectively in $\mathsf{A}_{\pm}$, $\mathsf{A}$ and $(g,Y)\mathsf{A}$. Suppose $\mathsf{W}_{\pm}$ and $\mathsf{W}$ be the corresponding vector subspaces parallel to $\mathsf{A}_{\pm}$ and $\mathsf{A}$. We simplify our notations and denote $\nu^{*}(\mathsf{W}_{\pm},\mathsf{W})$ by $\nu^{*}_{\pm,w}$, $\nu^{*}(\mathsf{W}_{\pm},g\mathsf{W})$ by $\nu^{*}_{\pm,gw}$ and $\nu^{*}(\mathsf{W}_{\pm},\mathsf{W}_{\mp})$ by $\nu^{*}_{\pm,\mp}$. Suppose $(\mathsf{W}_{\pm},\mathsf{W})=[h_{\pm,w}]\in\mathsf{G}/(\mathsf{P}^{+}\cap\mathsf{P}^{-})$. Then $(\mathsf{W}_{\pm},g\mathsf{W})=[gh_{\pm,w}]$. Hence, we use Lemma 4.1.3 and for any $h_{\pm,gw}\in\mathsf{G}$ which satisfy $(\mathsf{W}_{\pm},g\mathsf{W})=[h_{\pm,gw}]$ we obtain that
@@
Let $\mathsf{G}$ be a split semisimple Lie group, $\{g_{t}\}_{t\in(-1,1)}\subset\mathsf{G}$ be an one parameter family with $g_{0}=g$ loxodromic and $X=g^{-1}dg_{0}(1)$. Then
@@
$X=$ $\ \left.\frac{d}{dt}\right|_{t=0}g^{-1}g_{t}=\left.\frac{d}{dt}\right|_{t=0}h_{0}\exp(-\mathtt{Jd}(g))m^{-1}h_{0}^{-1}h_{t}m\exp(\mathtt{Jd}(g_{t}))h_{t}^{-1}$
$=$ $\ h_{0}\exp(-\mathtt{Jd}(g))m^{-1}h_{0}^{-1}dh_{0}(1)m\exp(\mathtt{Jd}(g))h_{0}^{-1}$
$+\ \left.\frac{d}{dt}\right|_{t=0}h_{0}\exp(-\mathtt{Jd}(g))\exp(\mathtt{Jd}(g_{t}))h_{0}^{-1}-dh_{0}(1)h_{0}^{-1}$
$=$ $\ g^{-1}dh_{0}(1)h_{0}^{-1}g+h_{0}\left(\left.\frac{d}{dt}\right|_{t=0}\mathtt{Jd}(g_{t})\right)h_{0}^{-1}-dh_{0}(1)h_{0}^{-1}.$
@@
$\mathtt{Ad}(h_{0}^{-1})[X]=\left.\frac{d}{dt}\right|_{t=0}\mathtt{Jd}(g_{t})+\mathtt{Ad}(h_{0}^{-1}gh_{0})[h_{0}^{-1}dh_{0}(1)]-h_{0}^{-1}dh_{0}(1).$
@@
Similar results relating Margulis invariants and Jordan projections has also been independently obtained by Andrés Sambarino and Kassel–Smilga [KSon].
@@ 4.3. Convexity of spectra
    $\beta(\gamma_{-},\eta_{-},\gamma_{+},\eta^{k}\gamma_{+})$ $+\beta(\gamma_{+},\eta_{+},\gamma_{-},\eta_{-k}\gamma_{-})$
@@ 5.1. Criteria for proper actions
Suppose $\{\gamma_{n}\}_{n\in\mathbb{N}}\subset\Gamma$ is a sequence such that
@@
We observe that a sequence $\{\gamma_{n}\}_{n\in\mathbb{N}}$ of elements of $\Gamma$ is divergent if and only if their translational length ${\ell(\gamma_{n})}$ diverges.
@@
As $X_{n}$ converge to $X$ and $\mathtt{M}_{(\rho,u)}(\gamma_{n})$ stays bounded, we conclude that the sequence $\{(\rho,u)(\gamma_{n})X_{n}\}_{n\in\mathbb{N}}$ also stays bounded. Therefore, the action of $(\rho,u)(\Gamma)$ on $\mathsf{V}$ is not proper. On the other way around, suppose the action of $(\rho,u)(\Gamma)$ on $\mathsf{V}$ is not proper. Hence, there exists a diverging sequence $\{\gamma_{n}\}_{n\in\mathbb{N}}$ of elements of $\Gamma$ and another sequence $\{X_{n}\}$ of points in $\mathsf{V}$ which converges to $X$ such that the sequence $\{(\rho,u)(\gamma_{n})X_{n}\}$ also converges to some point $Y$. Suppose $\lim_{n\to\infty}\gamma_{n}^{+}=a$ and $\lim_{n\to\infty}\gamma_{n}^{-}=b$. Let $P_{n}$ (resp. $Q_{n}$) be that unique point on $\mathsf{A}_{\gamma_{n}}$ such that
@@
Suppose $\sigma:\mathsf{U}\Gamma\to\mathcal{AV}$ is a neutralized section and $\tilde{\sigma}:\widetilde{\mathsf{U}_{0}\Gamma}\to\mathsf{V}$ be the map induced by the lift of $\sigma$. Then the following holds:
@@
$=\pi_{\gamma}[\tilde{\sigma}(\gamma_{+},\gamma_{-},\ell(\gamma))-\tilde{\sigma}(\gamma_{+},\gamma_{-},0)]$
@@
Moreover, as $0\in\mathtt{M}\text{-}\mathsf{Spec}(\rho,u)$, there exists a diverging sequence $\{\gamma_{n}\}_{n\in\mathbb{N}}$ such that
$\lim_{n\to\infty}\frac{\mathtt{M}_{\rho,u}(\gamma_{n})}{\ell(\gamma_{n})}=0.$
@@ 5.2. Deformation of free groups
Henceforth, we denote $\mathsf{SL}(n,\mathbb{R})$ by $\mathsf{SL}_{n}$ for notational ease. In this subsection we assume that $\Gamma$ is a non-abelian free group. Let $\iota:\mathsf{SL}_{2}\to\mathsf{SL}_{n}$ be the irreducible representation and let $d_{e}\iota:\mathfrak{sl}_{2}\to\mathfrak{sl}_{n}$ be the corresponding Lie algebra homomorphism. Then any injective representation $\rho:\Gamma\to\mathsf{SL}_{n}$ is called a Fuchsian representation if and only if $\rho=\iota\circ\tau$ for some $\tau:\Gamma\to\mathsf{SL}_{2}$. Moreover, any representation $\rho:\Gamma\to\mathsf{SL}_{n}$ which can be continuously deformed to obtain a Fuchsian representation is called a Hitchin representation. We denote the space of Hitchin representations by $\mathsf{Hit}(\Gamma,\mathsf{SL}_{n})$. Suppose $\rho$ be a Fuchsian representation and $\mathsf{T}_{[\rho]}\mathsf{Hit}(\Gamma,\mathsf{SL}_{n})$ be the tangent space at $[\rho]$.
@@
As $0\notin\mathtt{M}\text{-}\mathsf{Spec}(\rho,u)$, there exists some linear functional $\alpha:\mathfrak{sl}_{n}\to\mathbb{R}$ and $c>0$ such that $\alpha(\mathtt{M}_{\rho,u}(\gamma))\geq c\ell(\gamma)>0$ for all $\gamma\in\Gamma$. Let $\mathcal{A}_{\rho,u}:=(\rho,u)(\Gamma)\backslash(\widetilde{\mathsf{U}_{0}\Gamma}\times\mathfrak{sl}_{n})$. We use Theorem 3.8 of Hirsch-Pugh-Shub [HPS06] (see Section 6 of Bridgeman–Canary–Labourie–Sambarino [BCLS15] for more details) to obtain that there exists a family of Hölder continuous sections
@@
$\int_{\gamma}f_{\rho,u}=\frac{\mathtt{M}_{\rho,u}(\gamma)}{\ell(\gamma)}.$
It follows that for all non torsion elements $\gamma\in\Gamma$ we have
$\int_{\gamma}\alpha(f_{\rho,u})\geq c>0.$

### v3

@@ 2.1. Representation of Lie groups
We choose a connected component of $\mathfrak{a}\setminus\cup_{\alpha\in\Sigma}\ker(\alpha)$ and denote it by $\mathfrak{a}^{++}$. The closure of $\mathfrak{a}^{++}$ is denoted by $\mathfrak{a}^{+}$. Let $\mathsf{K}\subset\mathsf{G}$ (respectively $\mathsf{A}\subset\mathsf{G}$) be the connected subgroup whose Lie algebra is $\mathfrak{k}$ (respectively $\mathfrak{a}$) and let $\mathsf{A}^{+}:=\exp{(\mathfrak{a}^{+})}$. The Lie group $\mathsf{K}$ is a maximal compact subgroup of $\mathsf{G}$.
The subset of restricted roots which take positive values on $\mathfrak{a}^{+}$ is denoted by $\Sigma^{+}$. We note that $\Sigma=\Sigma^{+}\sqcup-\Sigma^{+}$ and define
@@
In order to guarantee existence of proper affine actions, henceforth, we only consider representations, $\mathtt{R}$, for which $\mathsf{V}^{0}$ is nontrivial.
Moreover, we assume that $\mathtt{R}$ admits an extreme, symmetric and generic vector and we fix one such vector $X_{0}$ in $\mathfrak{a}^{+}$. Such a representation $\mathtt{R}$ is called a non-swinging representation.
@@
Let $\Sigma^{+}(X_{0}):=\{\alpha\mid\alpha(X_{0})\geqslant 0\}$. We denote the vector subspace generated by $\mathfrak{g}^{0}$ and all $\mathfrak{g}^{\alpha}$ such that $\alpha\in\pm\Sigma^{+}(X_{0})$ by $\mathfrak{p}^{\pm,0}$.
Let $g\in\mathsf{G}$. We say $g$ is elliptic if $g$ is conjugate to some element in $\mathsf{K}$. We say $g$ is hyperbolic if $g$ is conjugate to some element in $\mathsf{A}$. We say $g$ is unipotent if $g$ is conjugate to some element in $\mathsf{N}$.
@@ 2.2. Anosov representations
The pair $(\mathsf{V}^{+,0},\mathsf{V}^{-,0})$ is a transverse pair of parabolic subspaces and for any pair of transverse parabolic spaces $(\mathsf{V}_{i},\mathsf{V}_{j})$, there exists $h_{i,j}\in\mathsf{G}$ such that $(\mathsf{V}_{i},\mathsf{V}_{j})=h_{i,j}(\mathsf{V}^{+,0},\mathsf{V}^{-,0})$.
If $\mathsf{V}_{i}$, $\mathsf{V}_{j}$ and $\mathsf{V}_{k}$ are three parabolic subspaces which are mutually transverse to each other and $h_{i,j},h_{i,k},h_{k,j}\in\mathsf{G}$ be such that
@@ 3.1. Neutral and neutralized sections
where the fiber over $p\in\widetilde{\mathsf{U}_{0}\Gamma}$ is denoted by $(p,\mathsf{V}^{+}_{p_{+}},\mathsf{V}^{0}_{p},\mathsf{V}^{-}_{p_{-}})$. We observe that $(p,\mathsf{V}^{+}_{p_{+}},\mathsf{V}^{0}_{p},\mathsf{V}^{-}_{p_{-}})=(p,g_{p}\mathsf{V}^{+},g_{p}\mathsf{V}^{0},g_{p}\mathsf{V}^{-})$ and $\mathsf{V}^{\pm}_{p_{\pm}}\oplus\mathsf{V}^{0}_{p}=\mathsf{V}^{\pm,0}_{p_{\pm}}$.
We observe that for any $Y\in\mathsf{V}^{0}$ we have $\nu_{Y}(hg)=h\nu_{Y}(g)$ for all $g,h\in\mathsf{G}$. Hence, it follows that $\rho(\gamma)\nu_{Y}\circ\xi(\gamma_{+},\gamma_{-})=\nu_{Y}\circ\xi(\gamma_{+},\gamma_{-})$. We abuse notation and denote $\nu_{Y}\circ\xi$ by $\nu_{Y}$. These maps induce the following sections which we again denote using the same notation:
@@ 3.2. Partial affine Anosov representations
We observe that $(g,Y)\mathsf{A}_{g,Y}=\mathsf{A}_{g,Y}$.
Suppose $\mathsf{A}_{1}$ and $\mathsf{A}_{2}$ are two such subspaces which are parallel to $\mathsf{V}^{0}_{g}$. Then $\mathsf{A}_{1}=\mathsf{A}_{2}+Z$ for some $Z\in\mathsf{V}$ and $gZ=Z$. Therefore, $Z\in\mathsf{V}^{0}_{g}$ and it implies that $\mathsf{A}_{1}=\mathsf{A}_{2}+Z=\mathsf{A}_{2}$, proving uniqueness. ∎
@@ 4.2. Margulis invariants
Suppose $X_{\pm},X,X_{g,Y}$ are any four points respectively in $\mathsf{A}_{\pm}$, $\mathsf{A}$ and $(g,Y)\mathsf{A}$. Suppose $\mathsf{W}_{\pm}$ and $\mathsf{W}$ be the corresponding vector subspaces parallel to $\mathsf{A}_{\pm}$ and $\mathsf{A}$. We simplify our notations and denote $\nu^{*}(\mathsf{W}_{\pm},\mathsf{W})$ by $\nu^{*}_{\pm,w}$, $\nu^{*}(\mathsf{W}_{\pm},g\mathsf{W})$ by $\nu^{*}_{\pm,gw}$ and $\nu^{*}(\mathsf{W}_{\pm},\mathsf{W}_{\mp})$ by $\nu^{*}_{\pm,\mp}$.
Suppose $(\mathsf{W}_{\pm},\mathsf{W})=[h_{\pm,w}]\in\mathsf{G}/(\mathsf{P}^{+}\cap\mathsf{P}^{-})$. Then $(\mathsf{W}_{\pm},g\mathsf{W})=[gh_{\pm,w}]$. Hence, we use Lemma 4.1.3 and for any $h_{\pm,gw}\in\mathsf{G}$ which satisfy $(\mathsf{W}_{\pm},g\mathsf{W})=[h_{\pm,gw}]$ we obtain that
@@
Let $\mathsf{G}$ be a split semisimple Lie group, $\{g_{t}\}_{t\in(-1,1)}\subset\mathsf{G}$ be an one parameter family with $g_{0}=g$ loxodromic and $X=dg_{0}(1)g^{-1}$. Then
@@
$X=$ $\ \left.\frac{d}{dt}\right|_{t=0}g_{t}g^{-1}=\left.\frac{d}{dt}\right|_{t=0}h_{t}\exp(\mathtt{Jd}(g_{t}))mh_{t}^{-1}h_{0}m^{-1}\exp(-\mathtt{Jd}(g))h_{0}^{-1}$
$=$ $\ dh_{0}(1)h_{0}^{-1}+\ \left.\frac{d}{dt}\right|_{t=0}h_{0}\exp(\mathtt{Jd}(g_{t}))\exp(-\mathtt{Jd}(g))h_{0}^{-1}$
$-h_{0}\exp(\mathtt{Jd}(g))mh_{0}^{-1}dh_{0}(1)h_{0}^{-1}h_{0}m^{-1}\exp(-\mathtt{Jd}(g))h_{0}^{-1}$
$=$ $\ dh_{0}(1)h_{0}^{-1}+h_{0}\left(\left.\frac{d}{dt}\right|_{t=0}\mathtt{Jd}(g_{t})\right)h_{0}^{-1}-gdh_{0}(1)h_{0}^{-1}g^{-1}.$
@@
$\mathtt{Ad}(h_{0}^{-1})[X]=\left.\frac{d}{dt}\right|_{t=0}\mathtt{Jd}(g_{t})+h_{0}^{-1}dh_{0}(1)-\mathtt{Ad}(h_{0}^{-1}gh_{0})[h_{0}^{-1}dh_{0}(1)].$
@@
Similar results relating Margulis invariants and Jordan projections has also been independently obtained by Andrés Sambarino [Sam24] and Kassel–Smilga [KSon].
@@ 4.3. Convexity of spectra
    $\beta(\gamma_{-},\eta_{-},\gamma_{+},\eta^{k}\gamma_{+})$ $+\beta(\gamma_{+},\eta_{+},\gamma_{-},\eta^{-k}\gamma_{-})$
@@ 5.1. Criteria for proper actions
Suppose $\{\gamma_{n}\}_{n\in\mathbb{N}}\subset\Gamma$ is a sequence such that the word length of $\gamma_{n}$ goes to infinity as $n$ goes to infinity and
@@
We observe that the translational length of a diverging sequence always diverges.
@@
As $X_{n}$ converge to $X$ and $\mathtt{M}_{(\rho,u)}(\gamma_{n})$ stays bounded, we conclude that the sequence $\{(\rho,u)(\gamma_{n})X_{n}\}_{n\in\mathbb{N}}$ also stays bounded. Therefore, the action of $(\rho,u)(\Gamma)$ on $\mathsf{V}$ is not proper.
On the other way around, suppose the action of $(\rho,u)(\Gamma)$ on $\mathsf{V}$ is not proper. We use Theorem 5.9 of Guichard–Wienhard [GW12] to deduce that there exists a diverging sequence $\{\gamma_{n}\}_{n\in\mathbb{N}}$ of elements of $\Gamma$ and another sequence $\{X_{n}\}_{n\in\mathbb{N}}$ of points in $\mathsf{V}$ which converges to $X$ such that the sequence $\{(\rho,u)(\gamma_{n})X_{n}\}_{n\in\mathbb{N}}$ also converges to some point $Y$ (for more details see the first two paragraphs of page 21 of Ghosh–Treib [GT17]). Suppose $\lim_{n\to\infty}\gamma_{n}^{+}=a$ and $\lim_{n\to\infty}\gamma_{n}^{-}=b$. Let $P_{n}$ (resp. $Q_{n}$) be that unique point on $\mathsf{A}_{\gamma_{n}}$ such that
@@
Suppose $t_{\gamma}$ is the period of the periodic orbit corresponding to $\gamma\in\Gamma$, $\sigma:\mathsf{U}\Gamma\to\mathcal{AV}$ is a neutralized section and $\tilde{\sigma}:\widetilde{\mathsf{U}_{0}\Gamma}\to\mathsf{V}$ be the map induced by the lift of $\sigma$. Then the following holds:
@@
$=\pi_{\gamma}[\tilde{\sigma}(\gamma_{+},\gamma_{-},t_{\gamma})-\tilde{\sigma}(\gamma_{+},\gamma_{-},0)]$
@@
Moreover, as $0\in\mathtt{M}\text{-}\mathsf{Spec}(\rho,u)$ and the metric $d$ on $\widetilde{\mathsf{U}_{0}\Gamma}$ is bi-Lipschitz to the product metric (see Corollary 8.3H of Gromov [Gro87]), there exists a diverging sequence $\{\gamma_{n}\}_{n\in\mathbb{N}}$ such that
$\lim_{n\to\infty}\frac{\mathtt{M}_{\rho,u}(\gamma_{n})}{t_{\gamma_{n}}}=0.$
@@ 5.2. Deformation of free groups
Henceforth, we denote $\mathsf{SL}(n,\mathbb{R})$ by $\mathsf{SL}_{n}$ for notational ease. In this subsection we assume that $\Gamma$ is a non-abelian free group. Let $\iota:\mathsf{SL}_{2}\to\mathsf{SL}_{n}$ be the irreducible representation and let $d_{e}\iota:\mathfrak{sl}_{2}\to\mathfrak{sl}_{n}$ be the corresponding Lie algebra homomorphism. Then any injective representation $\rho:\Gamma\to\mathsf{SL}_{n}$ is called a Fuchsian representation if and only if $\rho=\iota\circ\tau$ for some $\tau:\Gamma\to\mathsf{SL}_{2}$. Moreover, any representation $\rho:\Gamma\to\mathsf{SL}_{n}$ which can be continuously deformed to obtain a Fuchsian representation is called a Hitchin representation. We denote the space of Hitchin representations by $\mathsf{Hit}(\Gamma,\mathsf{SL}_{n})$. Suppose $\rho$ is a Fuchsian representation and $\mathsf{T}_{[\rho]}\mathsf{Hit}(\Gamma,\mathsf{SL}_{n})$ is the tangent space at $[\rho]$.
@@
As $0\notin\mathtt{M}\text{-}\mathsf{Spec}(\rho,u)$, there exists some linear functional $\alpha:\mathfrak{sl}_{n}\to\mathbb{R}$ and $c>0$ such that $\alpha(\mathtt{M}_{\rho,u}(\gamma))\geq c\ell(\gamma)>0$ for all $\gamma\in\Gamma$.
Let $\mathcal{A}_{\rho,u}:=(\rho,u)(\Gamma)\backslash(\widetilde{\mathsf{U}_{0}\Gamma}\times\mathfrak{sl}_{n})$. We use Theorem 3.8 of Hirsch-Pugh-Shub [HPS06] (see Section 6 of Bridgeman–Canary–Labourie–Sambarino [BCLS15] for more details) to obtain that there exists a family of Hölder continuous sections
@@
$\int_{\gamma}f_{\rho,u}=\frac{\mathtt{M}_{\rho,u}(\gamma)}{t_{\gamma}}.$
We use Corollary 8.3H of Gromov [Gro87] and obtain some $c_{0}>0$ such that for all non torsion elements $\gamma\in\Gamma$ we have
$\int_{\gamma}\alpha(f_{\rho,u})\geq c_{0}>0.$